# Eco-evolutionary dynamics of cooperative antimicrobial resistance in a population of fluctuating volume and size

### v1

@@ Abstract
Antimicrobial resistance to drugs (AMR), a global threat to human and animal health, is often regarded as resulting from a cooperative behaviour, leading to the coexistence of drug-resistant and drug-sensitive cells in large communities and static environments. However, microbial populations generally evolve in volatile environments, yielding sudden changes in the available amount of nutrients and toxins. These, together with demographic fluctuations (birth and death events), have been recognised to drastically alter the population size and strain coexistence, as well as the ability of one strain to fixate and take over the entire population. Motivated by the need to better understand the evolution of AMR, we study a population of time-varying size consisting of two competing strains, one drug-resistant and one drug-sensitive, subject to demographic and environmental variability. This is modelled by a binary carrying capacity randomly switching between mild and harsh environmental conditions, and driving the fluctuating volume (total amount of nutrients and antimicrobials at fixed concentration), and thus the size of the community (number of resistant and sensitive cells). We assume that AMR is a shared public good when the concentration of resistant cells exceeds a fixed concentration cooperation threshold, above which the sensitive strain has a growth advantage, whereas resistant cells dominate below it. Using computational means, and devising an analytical treatment (built on suitable quenched and annealed averaging procedures), we fully characterise the influence of fluctuations on the eco-evolutionary dynamics of AMR, and notably obtain specific strain fixation and long-lasting coexistence probabilities as a function of the environmental variation rate and cooperation threshold.
@@ 1 Introduction
The evolution of natural populations is often strongly influenced by environmental variability (EV), characterised by endlessly changing conditions, such as temperature, light, pH, toxins, and nutrients [CW81, Che94, VTL94, Che00, ESAH19, MHR21]. In particular, microbial populations generally evolve in volatile environments, subject to conditions fluctuating between harsh and mild. For instance, the abundance of nutrients in a community can undergo cycles of feast and famine and the amount of toxin can suddenly and radically change [VTL94, SK98, Pro09, HM20, MK18, WFM17, WFM18, AHNRM23]. Exogenous changes thus cause environmental fluctuations that shape the population evolution [KL05, AMvO08, LK14, ${}^{+}$11, ARTG21, RGOAC20], and in particular the ability of species to coexist [CW81, Che94, Che00, CCR14, KKS15, GM16, MS20, MSS21, Gri20, WS22, AHNRM23] and to act cooperatively [BBG07, Bro07, PW09, WGSV02, SG13, WFM18, HNARM23a]. Demographic noise (DN) stemming from random birth and death events is another source of fluctuations. These are negligible in large communities and strong in small communities, where DN can lead to the fixation of one species, when it takes over the entire population, and the extinction of the others [Ewe04, CK09, BM07, WSWP23]. The dynamics of the make-up and size of a population are often coupled [Rou79], resulting in its eco-evolutionary dynamics [PGH09, KA14, WFM18, TWAM20, TWMA23]. The joint effect of EV and DN is crucial in microbial communities where it can lead to population bottlenecks, resulting in colonies prone to fluctuations [WGSV02, PW09, BBG07, Bro07], and hence can greatly influence the eco-evolutionary dynamics of antimicrobial resistance (AMR) [${}^{+}$18, MB20, AHNRM23]. It is also worth mentioning the considerable recent efforts dedicated to researching the mechanisms underpinning the coexistence of competing species under various scenarios, see e.g. [${}^{+}$19, Pin19, SMM21, ${}^{+}$22]. Moreover, the influence of different kinds of variability (e.g. quenched disorder, heterogeneous rates) on species coexistence has notably been investigated in ecosystems exhibiting cyclic dominance, see e.g. [${}^{+}$14, DMPT18] and references therein.
@@
AMR, whose rise is a major societal threat [O’N16, Col22], can often be interpreted as resulting from a cooperative behaviour. This occurs when antimicrobial drugs are inactivated by an enzyme produced, at a metabolic cost, by resistant cells [Dav94, Wri05, ${}^{+}$13, VG14, BPF21, AHNRM23]. When the concentration of resistant microbes exceeds a fixed “concentration cooperation threshold” there are enough drug-inactivating enzymes for the protection against the toxin to be shared with sensitive cells at no metabolic cost. However, when the concentration of resistant cells is below the cooperation threshold, only resistant microbes benefit from the protection of the enzymes. In this scenario, AMR is mediated by drug-inactivating enzymes that act as public good only above the cooperation threshold. This yields the spread of the resistant strain below the threshold, where the drug limits the growth of (or kills) sensitive microbes, while these thrive above the cooperative threshold where the drug-inactivating enzyme concentration is high [Dav94, Wri05, ${}^{+}$13, AHNRM23]. In a large population and static environment, sensitive and resistant strains thus coexist [${}^{+}$13, VG14, ${}^{+}$15, BWB16]. However, this picture can be greatly altered by fluctuations arising in finite population subject to EV [WGSV02, BBG07, Bro07, PW09, ${}^{+}$12, SLKF12, ${}^{+}$18, AHNRM23].
@@
Motivated by the need to better understand the evolution of AMR, we study the influence of EV and DN on the eco-evolutionary dynamics of a population of time-varying size consisting of two competing species, an antimicrobial-resistant strain and another sensitive to antimicrobial drugs. Here, EV is modelled by a binary carrying capacity that randomly switches between states corresponding to mild and harsh conditions (high and low values, respectively), with the antimicrobial drug’s concentration kept constant. This choice of EV, that allows us to model population bottlenecks, is responsible for the fluctuating size of the community. This encompasses the time-variation of the number of cells of each strain, and the amount of nutrients and toxins in the community at fixed concentration (varying volume). Hence, this EV mimics naturalistic settings subject to sudden floods/drainage events that greatly alter their volume (such as in rivers, sewerage, and natural ponds) in the context of antimicrobial-polluted environments [GS15], as well as in microdroplet chemostat lab setups of time-varying volume [${}^{+}$13, ${}^{+}$20]. Complementarily, in references [HNARM23a, AHNRM23] EV was modelled by a binary carrying capacity switching at constant volume, i.e. a form of EV with time-varying concentration of nutrients and toxins. In what follows, we discuss the microbial behaviour under volume-fluctuating environments and, by combining analytical and computational means, we determine the environmental conditions for the long-lived coexistence of the species and the fixation properties of each strain. We determine the fixation-coexistence diagrams of the model, and find the nontrivial environmental conditions separating the phases of dominance/fixation and long-lived coexistence of the species. We rationalise our findings by devising an analytical approach built on suitable quenched and annealed averaging procedures, that allow us to capture the fixation-coexistence diagrams, and to obtain the conditional probability that resistant cells prevail as a function of the environmental variation rate and cooperation threshold.
@@ 2.1 Model
We consider a well-mixed population of time-varying size $N(t)=N_{R}(t)+N_{S}(t)$ consisting of $N_{R}$ resistant ($R$) microbes and $N_{S}$ sensitive ($S$) microbes, which compete for the same resources in the presence of an antimicrobial drug. The strain $R$ experiences a metabolic cost (reduced reproduction rate) to produce resistance enzymes that break down the drug and reduce its concentration in the community. Microbes of strain $S$ pay no metabolic cost associated to drug-inactivating enzymes (of which they are no producers), but their growth is hampered by the presence of the toxin. When the fraction of $R$ in the population is sufficiently high, they produce enough resistance enzymes to bring the drug below its minimum inhibitory concentration (MIC) [Dav94, Bro04, ${}^{+}$13, BWB16]. When this happens, the drug is ineffective and $S$ benefits from the enzymes protection despite not having to pay any cost for their production. The production of the resistance enzymes by $R$ making up a sufficiently high fraction of the community can hence be regarded as cooperative behaviour, and their protection against antimicrobials as a form of public good [${}^{+}$13, VG14, BPF21], see figure 1(a).
For concreteness, we assume a biostatic antimicrobial that only affects the growth rate of $S$  [^1] The case where the drug increases the death rate of the $S$ microbes corresponds to a biocidal antimicrobial, and is not directly considered here. This is a stringent limitation, since a given drug may act as a biocidal or biostatic antimicrobial depending on its concentration [SMM17].. The metabolic cost for $R$ to produce the enzymes is denoted by $s$ and the growth hindrance experienced by $S$ microbes in the presence of the drug (above the MIC) is denoted by $a$. In this setting, the population concentration of $R$, given by $x\equiv N_{R}/N$, determines whether the enzymes offer a public good protection to $S$. If $x_{\text{th}}$ is the fraction of $R$ necessary for the enzymes to share their protection across the community, $S$ is protected when $x\geq x_{\text{th}}$, whereas it is affected by the drug when $x<x_{\text{th}}$. Accordingly, the growth fitness $f_{R/S}$ of the $R$ and $S$ microbes are
@@
where $\theta[z]$ is the Heaviside step function, defined as $\theta[z]=1$ if ($z>0$) and $\theta[z]=0$ otherwise. The average population fitness is $\bar{f}=f_{R}N_{R}/N+f_{S}N_{S}/N$. It is worth noting that the concentration threshold $x_{\text{th}}$ is constant in this setting, while it varied in [HNARM23a] where the population was assumed to vary at constant volume.
The population thus evolves according to the multivariate birth-death process [Gar02, vK92, Ewe04] defined by
@@
with transition rates [WFM17, WFM18, TWAM20, SMM21]
@@
where the growth is limited by the logistic death rate $N/K$, and where $K$ is the carrying capacity, that is here assumed to be a time-fluctuating quantity, see below. This choice ensures that $N$ obeys the standard logistic dynamics in the mean field limit, see equation (7). In the expressions of $T^{+}_{R/S}$, without loss of generality and for mathematical convenience, we have normalised $f_{R/S}$ by the average fitness $\bar{f}$ [HNARM23a]. The growth rate of each strain is thus given by its fitness relative to the average population’s fitness, which is a common assumption for many biological and evolutionary processes [CW81, Ewe04, TH09]. This allows us to establish a neat relationship with the classical Moran process, the reference birth-death-like process used to model the evolution of idealised populations of constant total size [Mor62, Ewe04, AS06, BM07, CMF11].
@@ 2.2 Environmental fluctuations & master equation
Here, we model EV by allowing the carrying capacity $K$ to randomly switch between two values, where this represents changes in the volume (and therefore population size) of the system. This allows us to capture sudden extreme environmental changes, as often used in laboratory experiments [WGSV02, PW09, BBG07, Bro07, LK14, WFM17, HNARM23a, AMvO08, LK14, ARTG21, NLGS21], with a simple mathematical setup. Concretely, EV is driven by a dichotomous Markov noise (DMN) [Ben06, HL06, RDL11], with the transition
@@
occurring at rate $(1-\delta\xi)\nu$, where $\nu$ is the (average) switching rate and $\delta$ denotes the asymmetry in the DMN switching ($-1<\delta<1$, see below), and we refer to the value of $\xi$ as the environmental state ($\xi=1$ for the mild state, and $\xi=-1$ for the harsh). Here, the DMN is always at stationarity, implying that we have $\xi=\pm 1$ with probability $(1\pm\delta)/2$. The stationary DMN ensemble average is thus $\langle\xi(t)\rangle=\delta$ and its auto-covariance is $\langle\xi(t)\xi(t^{\prime})\rangle-\langle\xi(t)\rangle\langle\xi(t^{\prime})\rangle=(1-\delta^{2})e^{-2\nu|t-t^{\prime}|}$, which, up to a constant, coincides with the DMN auto-correlation. We note that $\nu$ is the average switching rate, and $1/(2\nu)$ is the DMN correlation time. Since we introduce the fluctuations in the environmental volume by changes in the carrying capacity, we model the binary switching carrying capacity by [WFM17, WFM18, WM20, TWAM20, HNARM23a, AHNRM23]
@@
Therefore, the master equation (ME) for the probability $P(N_{R},N_{S},\xi,t)$ that at time $t$ the population consists of $N_{R}$ and $N_{S}$ microbes and is in the environmental state $\xi$ is [Gar02]:
@@
Since $T^{\pm}_{R/S}=0$ when $N_{R/S}=0$, there is extinction of $R$ and fixation of $S$ ($N_{S}=N$), or fixation of $R$ ($N_{R}=N$) and extinction of $S$. Here, the fixation of one strain is therefore accompanied by the extinction of the other, and when this occurs the population composition no longer varies (even though its size continues to fluctuate) [^2] The final state of this is the absorbing state $N_{R}=N_{S}=0$, which corresponds to the eventual collapse and extinction of the population. This occurs after a time growing exponentially with the size of the system [SDF17, WFM17, WFM18, TWAM20]. Here, where $K(t)\gg 1$, this phenomenon is practically unobservable and will therefore not be considered.. The multivariate ME (2.2) can be simulated exactly using standard stochastic methods [Gil76, GB00, And07, HNARM23a], and encodes the eco-evolutionary dynamics of the model. In all our simulations, $\xi$ is at stationarity, and we thus initially have $\xi(0)=\pm 1$ and hence $K(0)=K_{\pm}$ with probability $(1\pm\delta)/2$, see appendix section A. For convenience and without loss of generality, the initial population size is chosen to coincide with the value of the carrying capacity at $t=0$, i.e. $N(0)=K(0)$. We have checked that this choice has no influence on the quasi-stationary distribution of $N$, and therefore on the discussion that follows.

### v2

@@ Abstract
Antimicrobial resistance to drugs (AMR), a global threat to human and animal health, is often regarded as resulting from a cooperative behaviour. Moreover, microbes generally evolve in volatile environments that, together with demographic fluctuations (birth and death events), drastically alter population size and strain survival. Motivated by the need to better understand the evolution of AMR, we study a population of time-varying size consisting of two competing strains, one drug-resistant and one drug-sensitive, subject to demographic and environmental variability. This is modelled by a binary carrying capacity randomly switching between mild and harsh environmental conditions, and driving the fluctuating volume (total amount of nutrients and antimicrobials at fixed concentration), and thus the size of the community (number of resistant and sensitive cells). We assume that AMR is a shared public good when the concentration of resistant cells exceeds a fixed concentration cooperation threshold, above which the sensitive strain has a growth advantage, whereas resistant cells dominate below it. Using computational means, and devising an analytical treatment (built on suitable quenched and annealed averaging procedures), we fully characterise the influence of fluctuations on the eco-evolutionary dynamics of AMR, and notably obtain specific strain fixation and long-lasting coexistence probabilities as a function of the environmental variation rate and cooperation threshold. We find that microbial strains tend to coexistence, but demographic fluctuations eventually lead to the extinction of resistant or sensitive cells for small or large values of the concentration cooperation threshold, respectively. This also holds for dynamic environments, whose specific properties determine the extinction timescale.
@@ 1 Introduction
The evolution of natural populations is often strongly influenced by environmental variability (EV), characterised by endlessly changing conditions, such as temperature, light, pH, toxins, and nutrients [1, 2, 3, 4, 5, 6]. In particular, microbial populations generally evolve in volatile environments, subject to conditions fluctuating between harsh and mild. For instance, the abundance of nutrients in a community can undergo cycles of feast and famine and the amount of toxin can suddenly and radically change [3, 7, 8, 9, 10, 11, 12, 13]. Exogenous changes thus cause environmental fluctuations that shape the population evolution [14, 15, 16, 17, 18, 19], and in particular the ability of species to coexist [1, 2, 4, 20, 21, 22, 23, 24, 25, 26, 13] and to act cooperatively [27, 28, 29, 30, 31, 12, 32]. Demographic noise (DN) stemming from random birth and death events is another source of fluctuations. These are negligible in large communities and strong in small communities, where DN can lead to the fixation of one species, when it takes over the entire population, and the extinction of the others [33, 34, 35, 36]. The dynamics of the make-up and size of a population are often coupled [37], resulting in its eco-evolutionary dynamics [38, 39, 12, 40, 41]. The joint effect of EV and DN is crucial in microbial communities where it can lead to population bottlenecks, resulting in colonies prone to fluctuations [30, 29, 27, 28], and hence can greatly influence the eco-evolutionary dynamics of antimicrobial resistance (AMR) [42, 43, 13]. It is also worth mentioning the considerable recent efforts dedicated to researching the mechanisms underpinning the coexistence of competing species under various scenarios, see e.g. [44, 45, 46, 47]. Moreover, the influence of different kinds of variability (e.g. quenched disorder, heterogeneous rates) on species coexistence has notably been investigated in ecosystems exhibiting cyclic dominance, see e.g. [48, 49] and references therein.
@@
AMR, whose rise is a major societal threat [50, 51], can often be interpreted as resulting from a cooperative behaviour. This occurs when antimicrobial drugs are inactivated by an enzyme produced, at a metabolic cost, by resistant cells [52, 53, 54, 55, 56, 13]. When the concentration of resistant microbes exceeds a fixed “concentration cooperation threshold” there are enough drug-inactivating enzymes for the protection against the toxin to be shared with sensitive cells at no metabolic cost. However, when the concentration of resistant cells is below the cooperation threshold, only resistant microbes benefit from the protection of the enzymes. In this scenario, AMR is mediated by drug-inactivating enzymes that act as public good only above the cooperation threshold. This yields the spread of the resistant strain below the threshold, where the drug limits the growth of (or kills) sensitive microbes, while these thrive above the cooperative threshold where the drug-inactivating enzyme concentration is high [52, 53, 54, 13]. In a large population and static environment, sensitive and resistant strains thus coexist [54, 55, 57, 58]. However, this picture can be greatly altered by fluctuations arising in finite population subject to EV [30, 27, 28, 29, 59, 60, 42, 13].
@@
Motivated by the need to better understand the evolution of AMR, we study the influence of EV and DN on the eco-evolutionary dynamics of a population of time-varying size consisting of two competing species, an antimicrobial-resistant strain and another sensitive to antimicrobial drugs. Here, EV is modelled by a binary carrying capacity that randomly switches between states corresponding to mild and harsh conditions (high and low values, respectively), with the antimicrobial drug’s concentration kept constant. This choice of EV, that allows us to model population bottlenecks, is responsible for the fluctuating size of the community. This encompasses the time-variation of the number of cells of each strain, and the amount of nutrients and toxins in the community at fixed concentration (varying volume). Hence, this EV mimics naturalistic settings subject to sudden floods/drainage events that greatly alter their volume (such as in rivers, sewerage, and natural ponds) in the context of antimicrobial-polluted environments [61], as well as in microdroplet chemostat lab setups of time-varying volume [62, 63]. Complementarily, in references [32, 13] EV was modelled by a binary carrying capacity switching at constant volume, i.e. a form of EV with time-varying concentration of nutrients and toxins. In what follows, we discuss the microbial behaviour under volume-fluctuating environments and, by combining analytical and computational means, we determine the environmental conditions for the long-lived coexistence of the species and the fixation properties of each strain. We determine the fixation-coexistence diagrams of the model, and find the nontrivial environmental conditions separating the phases of dominance/fixation and long-lived coexistence of the species. We rationalise our findings by devising an analytical approach built on suitable quenched and annealed averaging procedures, that allow us to capture the fixation-coexistence diagrams, and to obtain the conditional probability that resistant cells prevail as a function of the environmental variation rate and cooperation threshold.
@@ 2.1 Model
We consider a well-mixed population of time-varying size $N(t)=N_{R}(t)+N_{S}(t)$ consisting of $N_{R}$ resistant ($R$) microbes and $N_{S}$ sensitive ($S$) microbes, which compete for the same resources in the presence of a constant influx of antimicrobial drug. The strain $R$ experiences a fixed metabolic cost (reduced reproduction rate) to constantly produce resistance enzymes [^1] As opposed to the case where $R$ can regulate the production of resistance enzyme by quorum sensing [64, 65]. The influence of this mechanism on the evolution of AMR is not addressed in this work. that break down the drug in their immediate vicinity, protecting them and also reducing the overall drug concentration in the community. Microbes of strain $S$ pay no metabolic cost associated to drug-inactivating enzymes (of which they are no producers), but their growth is hampered by the presence of the toxin. When the fraction of $R$ in the population is sufficiently high, they produce enough resistance enzymes to bring the overall drug concentration below its minimum inhibitory concentration (MIC) [52, 66, 54, 58]. When this happens, the drug is ineffective everywhere and $S$ benefits from the enzymes protection despite not having to pay any cost for their production. The production of the resistance enzymes by $R$ making up a sufficiently high fraction of the community can hence be regarded as cooperative behaviour, and their protection against antimicrobials as a form of public good [54, 55, 56], see figure 1(a).
For concreteness, we assume a biostatic antimicrobial that only affects the growth rate of $S$ [^2] The case where the drug increases the death rate of the $S$ microbes corresponds to a biocidal antimicrobial, and is not directly considered here. This is a stringent limitation, since a given drug may act as a biocidal or biostatic antimicrobial depending on its concentration [67]. The metabolic cost for $R$ to produce the enzymes is denoted by $s$ and the growth hindrance experienced by $S$ microbes in the presence of the drug (above the MIC) is denoted by $a$. In this setting, the fraction of $R$ in the population, given by $x\equiv N_{R}/N$, determines whether the enzymes offer a public good protection to $S$. If $x_{\text{th}}$ is the fraction of $R$ necessary for the enzymes to share their protection across the community, $S$ is protected when $x\geq x_{\text{th}}$, whereas it is affected by the drug when $x<x_{\text{th}}$. Accordingly, the growth fitness $f_{R/S}$ of the $R$ and $S$ microbes are
@@
where $\theta[z]$ is the Heaviside step function, defined as $\theta[z]=1$ if ($z>0$) and $\theta[z]=0$ otherwise, $s$ is the resistance metabolic cost, and $a$ is the drug-driven growth hindrance [68, 69, 67, 70, 71]. The average population fitness is $\bar{f}=f_{R}N_{R}/N+f_{S}N_{S}/N$. It is worth noting that the concentration threshold $x_{\text{th}}$ is constant in this setting, while it varied in [32] where the population was assumed to vary at constant volume.
The population thus evolves according to the multivariate birth-death process [72, 73, 33] defined by
@@
with transition rates [11, 12, 40, 46]
@@
where the growth is limited by the logistic death rate $N/K$, and where $K$ is the carrying capacity, that is here assumed to be a time-fluctuating quantity, see below. This choice ensures that $N$ obeys the standard logistic dynamics in the mean field limit, see equation (7). In the expressions of $T^{+}_{R/S}$, without loss of generality and for mathematical convenience, we have normalised $f_{R/S}$ by the average fitness $\bar{f}$ [32]. The growth rate of each strain is thus given by its fitness relative to the average population’s fitness, which is a common assumption for many biological and evolutionary processes [1, 33, 74]. This allows us to establish a neat relationship with the classical Moran process, the reference birth-death-like process used to model the evolution of idealised populations of constant total size [75, 33, 76, 35, 77].
@@ 2.2 Environmental fluctuations & master equation
Here, we model EV by allowing the carrying capacity $K$ to randomly switch between two values, where this represents changes in the volume (and therefore population size) of the system. This allows us to capture sudden extreme environmental changes, as often used in laboratory experiments [30, 29, 27, 28, 16, 11, 32, 15, 16, 18, 78], with a simple mathematical setup. Concretely, EV is driven by a dichotomous Markov noise (DMN) [79, 80, 81], with the transition
@@
occurring at rate $(1-\delta\xi)\nu$, where $\nu$ is the (average) switching rate and $\delta$ denotes the asymmetry in the DMN switching ($-1<\delta<1$, see below), and we refer to the value of $\xi$ as the environmental state ($\xi=1$ for the mild state, and $\xi=-1$ for the harsh). Here, the DMN is always at stationarity, implying that we have $\xi=\pm 1$ with probability $(1\pm\delta)/2$. The stationary DMN ensemble average is thus $\langle\xi(t)\rangle=\delta$ and its auto-covariance is $\langle\xi(t)\xi(t^{\prime})\rangle-\langle\xi(t)\rangle\langle\xi(t^{\prime})\rangle=(1-\delta^{2})e^{-2\nu|t-t^{\prime}|}$, which, up to a constant, coincides with the DMN auto-correlation. We note that $\nu$ is the average switching rate, and $1/(2\nu)$ is the DMN correlation time. Since we introduce the fluctuations in the environmental volume by changes in the carrying capacity, we model the binary switching carrying capacity by [11, 12, 82, 40, 32, 13]
@@
Therefore, the master equation (ME) for the probability $P(N_{R},N_{S},\xi,t)$ that at time $t$ the population consists of $N_{R}$ and $N_{S}$ microbes and is in the environmental state $\xi$ is [72]:
@@
Since $T^{\pm}_{R/S}=0$ when $N_{R/S}=0$, there is extinction of $R$ and fixation of $S$ ($N_{S}=N$), or fixation of $R$ ($N_{R}=N$) and extinction of $S$. Here, the fixation of one strain is therefore accompanied by the extinction of the other, and when this occurs the population composition no longer varies (even though its size continues to fluctuate) [^3] The final state of this is the absorbing state $N_{R}=N_{S}=0$, which corresponds to the eventual collapse and extinction of the population. This occurs after a time growing exponentially with the size of the system [83, 11, 12, 40]. Here, where $K(t)\gg 1$, this phenomenon is practically unobservable and will therefore not be considered.. The multivariate ME (2.2) can be simulated exactly using standard stochastic methods [84, 85, 86, 32], and encodes the eco-evolutionary dynamics of the model. In all our simulations, $\xi$ is at stationarity, and we thus initially have $\xi(0)=\pm 1$ and hence $K(0)=K_{\pm}$ with probability $(1\pm\delta)/2$, see appendix section A. For convenience and without loss of generality, the initial population size is chosen to coincide with the value of the carrying capacity at $t=0$, i.e. $N(0)=K(0)$. We have checked that this choice has no influence on the quasi-stationary distribution of $N$, and therefore on the discussion that follows.